# Imitating Language via Scalable Inverse Reinforcement Learning

### v1

@@ A.1.1 IQLearn
For IQLearn, fine-tuning starts with a policy which is subsequently finetuned as a Q-function. In order to do so, we take the logits underlying the LLM softmax layer and continue training as Q-values. In entropy regularised RL, we can build on the equality between optimal policy $\pi^{*}$ and optimal Q-function $Q^{*}$ via $\pi^{*}(a|s)=\nicefrac{{1}}{{Z_{s}}}\exp{Q^{*}(s,a)}$ with normalization factor $Z_{s}=\sum_{a^{\prime}\in A}\exp{Q^{*}(s,a^{\prime})}$. We further obtain $V$ via $V(s)=\log\sum_{a\in A}\exp{Q(s,a)}$.
Due to the translation invariance of the softmax function, the there can be a considerable state-based offset between initial logits and final values. In other words, with $Q(s,a)=V(s)+A(s,a)$ only the advantage function $A$ has to be accurately reflected by the initial policy logits, while the state-based value $V$ can be arbitrarily inaccurate. In practice, we can add further KL regularization, or separate the representation of state and advantage function, to stabilize training during the additional identification of correct offsets but found it to be unnecessary for our experiments in comparison to recent work [15].
We handle terminal states by setting their values to zero in line with the original IQLearn paper [20]. Improvements such as learned values for the terminal states [2, 15] did not contribute to improved performance in this setting.
@@ A.1.2 Online IQLearn
We also implemented an online version of IQLearn, which makes use of additional (non-expert) samples. In contrast to IQLearn [21] which makes use of these examples to estimate the initial value, $E_{\rho_{0}}v(s)$ (which is equivalent between online and offline data when the prompt is the same as in our case), we integrate the additional samples to relax the distribution matching loss:
@@
This formulation strongly relates to the offline version of the algorithm but additionally applies the temporal difference based regularisation term on the additional non-expert transitions.
@@ A.3.1 Quality-Diversity Evaluation
In addition to the Self-BLEU metric, we further add plots for model entropy and add the ROUGE-LSUM results in Figure 9. Further performance metrics like MAUVE [43] and BertScore [66] can be of use in the future to further represent human judgement and preferences.
@@ A.3.3 Toy Experiments for Offline and Online IRL with Autoregressive Generation
The toy scenario displayed by the MDP in Figure 11 is used to represent a key difference between many classical control settings and autoregressive language generation. The agent starts in the left black state and has to reach the green on the right. In each of the bottom states the agent has 2 actions, move left or move up. The transition dynamics are noisy so that the agent executes the unwanted action 10% of the time. There are two variants of the MDP, one where the agent can return from the top states by learning to execute the right action and one where it cannot. The latter represents one aspect of concatenation dynamics, the agent cannot return to the exact same state after sampling the wrong action. In a way, it cannot correct its behaviour exactly.
When training offline and online variants IQLearn in this setting with demonstrations without mistakes (and their correction), we clearly see that the online version of IQLearn outperforms offline learning by over 11% in success rates, while in the setting without recovery, the difference is considerably smaller and results are within each others confidence bounds with more variance for the offline agent.
@@ A.3.4 Analyzing GAIL Stabilization
We empirically find GAIL overall more complex to tune and stabilize and aim to provide further insights here. Intuitively, control for language modelling differs from many classical applications via large discrete action space and terminations not being state but action conditioned. In other words, the agent can directly chose to terminate. Therefore the agent can learn to directly terminate if the discriminator provides negative rewards. If we provide positive rewards, tuning those rewards becomes a complex task as seen in Figure 12, where without additional MLE objective the GAIL agent often uses the maximum sample length with correlated loss of performance (visualized by the drop on ROUGE-1 values). Additional MLE training can help to both shape policy behaviour but further also provide more relevant data for the discriminator [50]. Especially in high dimensional action spaces it otherwise becomes challenging to obtain a useful reward signal from the discriminator as seen in prior work [59].

### v2

@@ A.1.1 IQLearn
For IQLearn, fine-tuning starts with a policy which is subsequently finetuned as a Q-function. In order to do so, we take the logits underlying the LLM softmax layer and continue training as Q-values. In entropy regularized RL, we can build on the equality between optimal policy $\pi^{*}$ and optimal Q-function $Q^{*}$ via $\pi^{*}(a|s)=\nicefrac{{1}}{{Z_{s}}}\exp{Q^{*}(s,a)}$ with normalization factor $Z_{s}=\sum_{a^{\prime}\in A}\exp{Q^{*}(s,a^{\prime})}$. We further obtain $V$ via $V(s)=\log\sum_{a\in A}\exp{Q(s,a)}$.
Due to the translation invariance of the softmax function, the there can be a considerable state-based offset between initial logits and final values. In other words, with $Q(s,a)=V(s)+A(s,a)$ only the advantage function $A$ has to be accurately reflected by the initial policy logits, while the state-based value $V$ can be arbitrarily inaccurate. In practice, we can add further KL regularization, or separate the representation of state and advantage function, to stabilize training during the additional identification of correct offsets but found it to be unnecessary for our experiments in comparison to recent work [Cundy and Ermon, 2024].
We handle terminal states by setting their values to zero in line with the original IQLearn paper [Garg et al., 2022a]. Improvements such as learned values for the terminal states [Al-Hafez et al., 2023, Cundy and Ermon, 2024] did not contribute to improved performance in this setting.
@@ A.1.2 Online IQLearn
We also implemented an online version of IQLearn, which makes use of additional (non-expert) samples. In contrast to IQLearn [Garg et al., 2022b] which makes use of these examples to estimate the initial value, $E_{\rho_{0}}v(s)$ (which is equivalent between online and offline data when the prompt is the same as in our case), we integrate the additional samples to relax the distribution matching loss:
@@
This formulation strongly relates to the offline version of the algorithm but additionally applies the temporal difference based regularization term on the additional non-expert transitions. A similar algorithm is derived in Appendix D.1 [Sikchi et al., 2023] with strong results on classical, continuous control domains of its state-action value function based version. It further connects to prior methods enabling the integration of sub-optimal data into imitation learning [Ma et al., 2022, Kim et al., 2022] with a useful overview on the broader field of related methods in [Sikchi et al., 2023].
@@ A.3.1 Quality-Diversity Evaluation
In addition to the Self-BLEU metric, we further add plots for model entropy and add the ROUGE-LSUM results in Figure 9. Further performance metrics like MAUVE [Pillutla et al., 2022] and BertScore [Zhang et al., 2019] can be of use in the future to further represent human judgement and preferences.
@@ A.3.3 Toy Experiments for Offline and Online IRL with Autoregressive Generation
The toy scenario displayed by the MDP in Figure 11 is used to represent a key difference between many classical control settings and autoregressive language generation. The agent starts in the left black state and has to reach the green on the right. In each of the bottom states the agent has 2 actions, move left or move up. The transition dynamics are noisy so that the agent executes the unwanted action 10% of the time. There are two variants of the MDP, one where the agent can return from the top states by learning to execute the right action and one where it cannot. The latter represents one aspect of concatenation dynamics, the agent cannot return to the exact same state after sampling the wrong action. In a way, it cannot correct its behavior exactly.
When training offline and online variants IQLearn in this setting with demonstrations without mistakes (and their correction), the online version of IQLearn outperforms offline learning by over 11% in success rates, while in the setting without recovery, the difference is considerably smaller and results are within each others confidence bounds with more variance for the offline agent.
@@ A.3.4 Analyzing GAIL Stabilization
We empirically find GAIL overall more complex to tune and stabilize and aim to provide further insights here. Intuitively, control for language modelling differs from many classical applications via large discrete action space and terminations not being state but action conditioned. In other words, the agent can directly chose to terminate. Therefore the agent can learn to directly terminate if the discriminator provides negative rewards. If we provide positive rewards, tuning those rewards becomes a complex task as seen in Figure 12, where without additional MLE objective the GAIL agent often uses the maximum sample length with correlated loss of performance (visualized by the drop on ROUGE-1 values). Additional MLE training can help to both shape policy behavior but further also provide more relevant data for the discriminator [Ren et al., 2024]. Especially in high dimensional action spaces it otherwise becomes challenging to obtain a useful reward signal from the discriminator as seen in prior work [Wu et al., 2021].